OpenStudy (anonymous):

In March, Isaiah and Freddie had the same amount of money in their savings accounts. In April, Isaiah deposited $210 into his account. Freddie increased the money in his account by 15%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in March?

OpenStudy (texaschic101):

x = money in the accounts in March x + 210 = x + 0.15x 210 = 0.15x 210/0.15 = x 1400 = x check.. Isaah -- 1400 + 210 = 1610 Freddie -- 1400 + 0.15(1400) = 1400 + 210 = 1610 March accounts = $ 1400